OpenStudy (anonymous):

Find limit f(x)=The root of (x+3) - 1 / x+2 when x is heading towards -2

OpenStudy (ipwnbunnies):

Alright, so you understand when you plugged in -2 for x, you got the form 0/0. correct?

OpenStudy (anonymous):

yea

OpenStudy (ipwnbunnies):

Usually when you have square roots in a limit problem, you'll have to multiply the fraction by the conjugate of the square root term.

OpenStudy (ipwnbunnies):

For example, the conjugate of 'sqrt(x) - 2' is 'sqrt(x) + 2'

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2} \sqrt{x+3} +1 \div x+2\]

OpenStudy (anonymous):

yea

OpenStudy (ipwnbunnies):

Right, that's the function, we have a 0/0 form. Multiply both the numerator and the denominator by the conjugate of the square root term.

OpenStudy (anonymous):

u mean we multiple it with \[\sqrt{x+3} +1\div \sqrt{x+3} +1\]

OpenStudy (ipwnbunnies):

\[f(x) = \frac{ \sqrt(x+3) - 1 }{ x+2 } * \frac{\sqrt(x+3) +1}{\sqrt(x+3) + 1}\]

OpenStudy (ipwnbunnies):

Oh yeah, you got it. Sorry, I didn't read your post right at first glance lol.

OpenStudy (ipwnbunnies):

Hint: You don't have to expand the denominator usually.

OpenStudy (anonymous):

So i get this eventually \[\frac{ 1 }{ \sqrt{x+3} +1 }\] right ?

OpenStudy (ipwnbunnies):

Err, I'm not sure. Did you simplify the product?

OpenStudy (anonymous):

and eventually it becomes \[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Yea

OpenStudy (ipwnbunnies):

\[\lim_{x \rightarrow -2} \frac{ (x+3) - 1 }{ (x+2)(\sqrt(x+3) +1)}\]

OpenStudy (ipwnbunnies):

It shouldn't be reduced to 1/2 because of the (x+2) binomial in the denominator.

OpenStudy (anonymous):

I.k But x+3-1 = (x+2)

OpenStudy (anonymous):

we take out the x+2 that leavs us with \[\frac{ 1 }{ \sqrt{x+3}+1 }\]

OpenStudy (ipwnbunnies):

OH! Lol, correct, totally did not see that. It's early in the morning. So very sorry. Yes, now the problem has become the limit as x approaches -2 of the new function!

OpenStudy (ipwnbunnies):

And yes, it'll be 1/2.

OpenStudy (anonymous):

Hahahaha N.p Thanks btw

OpenStudy (ipwnbunnies):